A 350 mL sample of gas is collected over water at 30°C and at a total pressure of 640 mm Hg. If the vapor pressure of water at 30°C is 31.8 torr, what is the volume of the gas sample at standard pressure?


1
Expert's answer
2021-03-16T07:49:51-0400

31.0torr = 31.0 mmHg

Partial pressure of the gas = (640 - 31.0)mmHg

= 609mmHg

Changing pressure to atm = "=\dfrac{609mmHg}{760mmHg/atm}=0.8atm"


Use the formula;


"P1V1 =P2V2" (at constant temperature)


P1 = 0.8atm

P2=1atm (standard pressure)

V1=350mL

V2=?


"V2=\dfrac{P1V1}{P2}=\dfrac{0.8atm x 350mL}{1atm} =280mL"